OpenStudy (swagster23780):

Which ratio is equivalent to 14/18? A.1/2 B.8/14 C.1/3 D.2/7

OpenStudy (theprofessor35):

Multiply or divide

OpenStudy (theprofessor35):

14/2 over 18/2 =

OpenStudy (mathwizzard3):

simplify

OpenStudy (theprofessor35):

Well, actually, non of these

OpenStudy (theprofessor35):

Non of these are equal to 14/18

rebeccaxhawaii (rebeccaxhawaii):

none of these are correct

rebeccaxhawaii (rebeccaxhawaii):

14/18=0.777777

OpenStudy (mathwizzard3):

exactly

OpenStudy (theprofessor35):

Are you sure you copied the ratio correctly?

OpenStudy (swagster23780):

Yes

OpenStudy (swagster23780):

Sorry 14/28

OpenStudy (theprofessor35):

Ohhhh,

OpenStudy (theprofessor35):

14/28 14/2 = 7 28/2 = 14

OpenStudy (mathwizzard3):

Oh!

OpenStudy (theprofessor35):

7/14 is teh same as 1/2

OpenStudy (mathwizzard3):

yes 1/2

OpenStudy (theprofessor35):

the answer would be.....

OpenStudy (theprofessor35):

1/2! :))))